# Achronal localization and representation of the causal logic from conserved current, application to massive scalar boson

### v2

@@ Abstract
1 Introduction
Achronal localization. There are plain physical grounds why in order to comply with causality localization of a relativistic quantum mechanical system has to occur in all achronal regions of spacetime. This means that to every achronal Borel set $\Delta\subset\mathbb{R}^{4}$ there is assigned a nonnegative bounded operator $T(\Delta)$ acting on the Hilbert space of states with the expectation value $\langle\phi,T(\Delta)\phi\rangle$ being the probability of localization in $\Delta$ of the system in the state $\phi$. As argued further in [7] the probabilities of localization regarding achronally separated regions add up and yield $1$ for every maximal achronal region. In summary $T$ is a positive operator valued normalized measure (POVM) on every maximal achronal set. Moreover, by relativistic symmetry, $T$ is Poincaré covariant with respect to the unitary kinematical transformations of the states.
       Localization on spacelike Cauchy surfaces. Just in order to meet the requirements of causal localization [12] extended localization from flat spacelike regions, as up to then commonly considered, to proper spacetime regions studying thoroughly POVM on spacelike smooth Cauchy surfaces. The POVM are coherent in the sense that the localization operators assigned to a region contained in the intersection of two different spacelike smooth Cauchy surfaces coincide. An important result is that the localization considered in [12] satisfies a rather general causality requirement. The considerations actually extend to $C^{1}$ Cauchy surfaces. Here we do the final step to achronal localization described above.
       The construction of an achronal localization goes along the lines of the construction of the Cauchy localization in [12]. However there are to overcome some technical difficulties due to the fact that the Lipschitz continuous function, with its graph being the achronal set under consideration, in general is not $C^{1}$.
       Divergence theorem. For this purpose we prove the divergence theorem for open bounded subsets of $\mathbb{R}^{n}$ with almost Lipschitz boundary such that the boundary has finite $(n-1)$-dimensional Hausdorff measure and the irregular points of the boundary are contained in a compact set of zero $(n-1)$-dimensional Minkowski content. This extension of the divergence theorem (3) is new. In view of applications we add some remarks about how verify the assumptions of the theorem.
       In corollary (4) the result of (4) is extended to all maximal achronal sets. (4) is proven under an additional boundedness assumption. Regarding the application to the massive scalar boson this assumption turns out to be almost irrelevant.
       Application to the massive scalar boson. The localizations of the massive scalar boson in Euclidean space (i.e., the Euclidean covariant positive operator valued normalized measures in $\mathbb{R}^{3}$, called POL in [7]) are determined by the integrals of a positive definite rotational invariant probability density $J_{0}$ over the regions of localization [7, sec. 6]. $J_{0}$ is the zeroth component of a covariant conserved current $\mathfrak{J}$ if and and only if the kernel is causal (6). This result is by [14]. Causal kernels have been studied in [14], [15], [7]. Under the almost irrelevant condition that the causal kernel is $C^{4}$ it is easy to verify that $\mathfrak{J}$ satisfies the assumption in (5) thus giving rise to a covariant achronal localization of the massive scalar boson.
       Analogously one obtains a covariant family of covariant achronal localizations related to the stress energy tensor of the massive scalar boson [18].
       Representation of the causal logic. Every (covariant) achronal localization determines uniquely a (covariant) representation of the causal logic and vice versa [7, (21)]. Obviously this one-to-one correspondence emphasizes further the relevance of achronal localization.
       Hence for the first time a covariant representation of the causal logic for an elementary relativistic quantum mechanical system is achieved. Also a covariant family of representation of the causal logic is derived corresponding to the above mentioned family of achronal localizations related to the stress energy tensor.
@@ 2 Notations and notions
Vectors in $\mathbb{R}^{4}$ are denoted by $\mathfrak{x}=(x_{0},x)$ with $x:=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. Let $\varpi:\mathbb{R}^{4}\to\mathbb{R}^{3}$ denote the projection $\varpi(\mathfrak{x}):=x$. Representing Minkowski spacetime by $\mathbb{R}^{4}$ the Minkowski product of $\mathfrak{a},\mathfrak{a}^{\prime}\in\mathbb{R}^{4}$ is given by $\mathfrak{a}\cdot\mathfrak{a}^{\prime}:=a_{0}a_{0}^{\prime}-aa^{\prime}$, where for vectors $a,a^{\prime}$ in $\mathbb{R}^{3}$ the scalar product $a_{1}a^{\prime}_{1}+a_{2}a^{\prime}_{2}+a_{3}a^{\prime}_{3}$ is denoted by $aa^{\prime}$. Often we use the notation $\mathfrak{a}^{\cdot 2}:=\mathfrak{a}\cdot\mathfrak{a}$.
       $\tilde{\mathcal{P}}=ISL(2,\mathbb{C})$ is the universal covering group of the Poincaré group. It acts on $\mathbb{R}^{4}$ as
@@
       The fourvector $\mathfrak{z}\in\mathbb{R}^{4}\setminus\{0\}$ is called timelike, lightlike, causal if $|z_{0}|>|z|$, $|z_{0}|=|z|$, $|z_{0}|\geq|z|$, respectively. It is future-directed if $z_{0}>0$.
@@
       A set $\Delta\subset\mathbb{R}^{4}$ is said to be spacelike if $|x_{0}-y_{0}|<|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$, $\mathfrak{x}\neq\mathfrak{y}$.
@@
       A set $\Delta\subset\mathbb{R}^{4}$ is said to be achronal if $|x_{0}-y_{0}|\leq|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$. By definition $\Delta$ is maximal achronal if $\Delta$ in not properly contained in an achronal set. An achronal set is maximal achronal if and only if it meets every timelike line. Note also that a maximal achronal set is always closed.
@@ 3 Divergence theorem on open sets with almost Lipschitz boundary
The following is inspired by Maggi [17, Remark 9.5, Theorem 9.6]. As to the notation, for $x\in\mathbb{R}^{n}$, $z\in\mathbb{R}^{n-1}$, $s>0$ put $x^{\prime}:=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$, and let $B_{s}(x):=\{y\in\mathbb{R}^{n}:|y-x|<s\}$, $C(x,s):=\{y\in\mathbb{R}^{n}:|y^{\prime}-x^{\prime}|<s,|y_{n}-x_{n}|<s\}$, $D(z,s):=\{y\in\mathbb{R}^{n-1}:|y-z|<s\}$. Given an outer measure $\omega$ on $\mathbb{R}^{n}$ and $X\subset\mathbb{R}^{n}$, the trace (restriction) of $\omega$ on $X$ is the outer measure $\omega|X$ on $\mathbb{R}^{n}$ given by $\omega|X(F):=\omega(F\cap X)$. $\mathcal{H}^{n-1}$ denotes the $n-1$-dimensional Hausdorff measure on $\mathbb{R}^{n}$.
@@
       $\mathcal{H}^{n-1}|C(x,s)\cap M=\mathcal{H}^{n-1}|\operatorname{graph}\,u|_{D}$ is a Radon measure due to the area formula ([17, Theorem 9.1]) . Moreover, one has $\mathcal{H}^{n-1}(\operatorname{graph}\,u|_{D(x^{\prime},s)\setminus G})=\int_{D(x^{\prime},s)\setminus G}|(\nabla u(z),-1)|\operatorname{d}z=0$. ∎
The following preparatory result (3) comprises by the case $M_{0}=\emptyset$ the divergence theorem on bounded open sets with Lipschitz boundary (see also e.g. [17, Remark 9.5] and [1, A8.8]).
@@
Proof. (a) Assume first $\varphi\in C_{c}^{1}(C(x,s))$. By (3) the surface integral is well-defined. Following step one of the proof of [17, Theorem 9.3] the result follows due to $C(x,s)\cap\partial E=C(x,s)\cap M$ by definition (3)(b). Note that the weak gradient $\nabla f_{\delta}$ equals $\frac{1}{2\delta}(-\nabla u(z),1)$ at $z\in G$.
       (b) Now the vector field $\nu_{E}$ is constructed. Applying the result in (a) to all $\varphi\in C_{c}^{1}(C(x,s)\cap C(\bar{x},\bar{s}))$ shows that $\nu$ and $\bar{\nu}$ coincide a.e. on their common domain. There are countably many $C(x_{m},s_{m})$ from the localization in (3), which cover $M$. (Indeed, $M_{0}$ is closed and hence $G_{\delta}$. Since $\partial E$ is compact it follows that $M=\partial E\setminus M_{0}$ is $\sigma$-compact.) Thus we may compose from the corresponding $\nu_{m}$ by means of [17, Theorem 9.1] an $\mathcal{H}^{n-1}$-a.e. determined unit vector field $\nu_{E}$ on $\partial E$ which locally coincides a.e. with $\nu$ from (3).
@@
Indeed, the first claim holds as $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})<\infty$ and $A_{\delta}\downarrow_{\delta}M_{0}$. The very definition of the Minkowski content [13, 3.2.27] implies the second claim.
@@
       According to [17, Theorem 8.1], $\mathcal{H}^{n-1}(D)<\infty$ if $D$ is the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable set $C$ of finite Lebesgue measure. If $C$ is a Lebesgue null set then $\mathcal{H}^{n-1}(D)=0$.
@@
       According to [13, Theorem 3.2.39], $\mathcal{H}^{n-1}(D)=\mathcal{M}^{n-1}(D)=0$ holds, if $D\subset\mathbb{R}^{n}$ is closed and if there is a Lipschitz function $h:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ mapping some bounded $C\subset\mathbb{R}^{n-1}$ onto $D$ or, equivalently, if $h$ is locally Lipschitz defined on $\overline{C}$ with $h(C)=D$. The latter holds because $\overline{C}$ is compact.

### v4

# Achronal localization and representation of the causal logic from a conserved current, application to the massive scalar boson

@@ 1 Notations and notions
Minkoski spacetime. Vectors in $\mathbb{R}^{4}$ are denoted by $\mathfrak{x}=(x_{0},x)$ with $x:=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. Let $\varpi:\mathbb{R}^{4}\to\mathbb{R}^{3}$ denote the projection $\varpi(\mathfrak{x}):=x$. Representing Minkowski spacetime by $\mathbb{R}^{4}$ the Minkowski product of $\mathfrak{a},\mathfrak{a}^{\prime}\in\mathbb{R}^{4}$ is given by $\mathfrak{a}\cdot\mathfrak{a}^{\prime}:=a_{0}a_{0}^{\prime}-aa^{\prime}$, where for vectors $a,a^{\prime}$ in $\mathbb{R}^{3}$ the scalar product $a_{1}a^{\prime}_{1}+a_{2}a^{\prime}_{2}+a_{3}a^{\prime}_{3}$ is denoted by $aa^{\prime}$. Often we use the notation $\mathfrak{a}^{\cdot 2}:=\mathfrak{a}\cdot\mathfrak{a}$.
Poincaré group. $\tilde{\mathcal{P}}=ISL(2,\mathbb{C})$ is the universal covering group of the proper orthocronous Poincaré group. It acts on $\mathbb{R}^{4}$ as
@@
Spacetime relations. The fourvector $\mathfrak{z}\in\mathbb{R}^{4}\setminus\{0\}$ is called timelike, lightlike, causal if $|z_{0}|>|z|$, $|z_{0}|=|z|$, $|z_{0}|\geq|z|$, respectively. It is future-directed if $z_{0}>0$.
@@
       A set $\Delta\subset\mathbb{R}^{4}$ is said to be spacelike if different points of $\Delta$ are spacelike separated, i.e., $|x_{0}-y_{0}|<|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$, $\mathfrak{x}\neq\mathfrak{y}$.
Achronal sets. A set $\Delta\subset\mathbb{R}^{4}$ is said to be achronal if the points of $\Delta$ are achronal separated, i.e., $|x_{0}-y_{0}|\leq|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$.
@@
       Clearly every spacelike set is achronal. By definition $\Delta$ is maximal achronal if $\Delta$ is not properly contained in an achronal set. An achronal set is maximal achronal if and only if it meets every timelike line.
       Cauchy surfaces in Minkowski space are maximal achronal sets, which meet all lightlike lines. The converse holds also true [8, Moretti (36)]. The mass shell is an example of a spacelike maximal achronal set, which is not a Cauchy surface. Being maximal it is not even contained in a Cauchy surface.
       Note also that every achronal set is contained in a maximal one and that a maximal achronal set is always closed.
       Most important is the fact that a set $\Delta\subset\mathbb{R}^{4}$ is achronal if and only if it is the graph of a $1$-Lipschitz function $\tau:\varpi(\Delta)\to\mathbb{R}$. It is maximal achronal iff $\varpi(\Delta)=\mathbb{R}^{3}$.
       Let the maximal achronal set $\Lambda=\{\big{(}\tau(x),x\big{)}:x\in\mathbb{R}^{3}\}$ be endowed with the induced topology from $\mathbb{R}^{4}$. Then $j:\mathbb{R}^{3}\to\Lambda$, $j(x):=(\tau(x),x)$ is a homeomorphism, $\{\mathfrak{x}\in\mathbb{R}^{4}:x_{0}<\tau(x)\}$ and $\{\mathfrak{x}\in\mathbb{R}^{4}:x_{0}>\tau(x)\}$ are open with boundary $\Lambda$, and $\Lambda$ is path-connected. In particular $\Lambda$ is a topological hypersurface. Clearly $\Lambda\subset\mathbb{R}^{4}$ has no interior points.
       For more details on achronal sets see [8, sec. 2.2].
2 Introduction
Achronal localization. There are plain physical grounds why in order to comply with causality, localization of a relativistic quantum mechanical system has to occur in all achronal regions of spacetime. This means that to every achronal Borel set $\Delta\subset\mathbb{R}^{4}$ (see sec. 1, Achronal sets) there is assigned an effect: a nonnegative operator $T(\Delta)$, bounded by $I$, acting on the Hilbert space of states. The expectation value $\langle\phi,T(\Delta)\phi\rangle$ is the probability of localization in $\Delta$ of the system in the state $\phi$. As argued further in [8] the probabilities of localization regarding achronally separated regions countably add up and yield $1$ for every maximal achronal region. In summary $T$ is a normalized positive operator valued measure (POVM) on every maximal achronal set. Moreover, by relativistic symmetry, $T$ is Poincaré covariant with respect to the unitary representation of the Poincaré group corresponding to kinematical transformations of the states.
       Achronal localization is causal. It satisfies the so-called Einstein causality in the most direct interpretation, which requires from localization that the probability of localization in a region of influence cannot be less than that in the region of actual localization. Details about this causality condition CC and achronal localization are furnished in sec. 5.
       Heuristically one may imagine a massive free particle to be represented by a timelike straight world line. The particle is considered to be localized in a spacetime region if its world line crosses the region. This simple heuristic picture of localizability implies rather stringent achronal localization [8, sec. 3.3]. Since the maximal achronal sets are those sets, which are met by every timelike straight line just once, one infers that a massive particle is localized in every maximal achronal region. Hence normalization follows. The present results on the localization of the massive scalar boson evidence the usefulness of this heuristical reasoning.
       The question is how to realize an apparatus, which ascertains the presence of a particle in an achronal spacetime region $\Delta$. From a physical point of view it is reasonable to assume the region to be spacelike piecewise flat. Presume that localization in $3D$ Euclidean space is feasible. This implies by relativistic symmetry that for every flat piece $\Gamma$ there is an apparatus suited to ascertain the presence of the particle in that region $\Gamma$. Being spacelike separated all these apparatus constitute one apparatus able to ascertain the presence of the particle in $\Delta$.
       Localization on spacelike Cauchy surfaces. Just in order to meet the requirements of causal localization [13] extended localization from flat spacelike regions, as up to then commonly considered, to proper spacetime regions studying thoroughly POVM on spacelike smooth Cauchy surfaces (which are a special case of maximal achronal sets). The POVM are coherent in the sense that the localization operators assigned to a region contained in the intersection of two different spacelike smooth Cauchy surfaces coincide. An important result is that the localization considered in [13] satisfies a rather general causality requirement. The results in [13] were achieved by a suitable use of the divergence theorem for volume forms. It was done by taking advantage of relatively recent advanced results in Lorentzian geometry concerning the extension of acausal manifolds with boundary to spacelike smooth Cauchy surfaces. These achievements actually straightforwardly extend to spacelike $C^{1}$ Cauchy surfaces.
Here we do the final step towards a fully-fledged achronal localization notion as described above.
       The construction of an achronal localization goes along the lines of the construction of the Cauchy localization in [13]. However there are some technical difficulties, which arise due to the facts (i) that a maximal achronal set in general is not a Cauchy surface as it is not met by every lightlike line and (ii) that an achronal set in general is not $C^{1}$, which means that it is not the graph of a $C^{1}$ function. Achronal sets are representable by Lipschitz functions, this only regularity condition should be therefore assumed.
       Divergence theorem. The divergence theorem plays a crucial role in the construction of concrete achronal localization notions. In order to overcome the difficulties pointed out above, we prove the divergence theorem for open bounded subsets of $\mathbb{R}^{n}$ with almost Lipschitz boundary such that the boundary has finite $(n-1)$-dimensional Hausdorff measure and the irregular points of the boundary are contained in a compact set of zero $(n-1)$-dimensional Minkowski content. This extension (3) of the divergence theorem is new. In view of applications we add some remarks about how to verify the assumptions of the theorem. At the beginning of sec. 4 it is expounded how (3) applies in showing the crucial technical result (4).
       The result of (4) is extended to all maximal achronal sets containing the origin under the assumption on the decay (5)(b) of the probability current. Due to this decay the extension is achieved simply by flatten the maximal achronal sets becoming $\gamma$-achronal for $\gamma<1$. The assumed decay is determined by the free relativistic time evolution of a massive particle. Regarding the application to the massive scalar boson this assumption turns out to be a technicality satisfied for all relevant currents.
       Application to the massive scalar boson. The localizations of the elementary particle being the massive scalar boson in Euclidean space (i.e., the Euclidean covariant positive operator valued normalized measures in $\mathbb{R}^{3}$, called POL in [8]) are determined by the integrals of a positive definite rotational invariant probability density $J_{0}$ over the regions of localization [8, sec. 6]. $J_{0}$ is the zeroth component of a covariant conserved current $\mathfrak{J}$ if and and only if the kernel is causal (6). This result is by [15]. Causal kernels have been studied in [15], [19], [8]. Under the physically irrelevant condition that the causal kernel is $C^{4}$ one verifies that $\mathfrak{J}$ satisfies the assumption in (5) thus giving rise to a covariant achronal localization of the massive scalar boson.
       Analogously one obtains a covariant family of achronal localizations related to the stress energy tensor of the massive scalar boson [23].
       Representation of the causal logic. Every (covariant) achronal localization determines uniquely a (covariant) representation of the causal logic and vice versa [8, (21)]. Obviously this one-to-one correspondence emphasizes further the relevance of achronal localization.
       Hence for the first time a covariant representation of the causal logic for an elementary relativistic quantum mechanical system is achieved. Also a covariant family of representations of the causal logic is derived corresponding to the above mentioned family of achronal localizations related to the stress energy tensor.
@@ 3 Divergence theorem on open sets with almost Lipschitz boundary
The following is inspired by Maggi [21, Remark 9.5, Theorem 9.6]. Repeatedly one refers to the well-known Area Formula in geometric measure theory first proved by H. Federer, see [14, 3.2., in particular 3.2.20], citing the easier accessible [21, Theorem 9.1]. Part (a) of the proof of the preparatory result (3) is the standard step familiar from the proof of the classical Gauß’s theorem. See e.g. [1, proof of (A8-16)] or the more elementary Step one of the proof of [21, Theorem 9.3].
As to the notation, for $x\in\mathbb{R}^{n}$, $z\in\mathbb{R}^{n-1}$, $s>0$ put $x^{\prime}:=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$, and let $B_{s}(x):=\{y\in\mathbb{R}^{n}:|y-x|<s\}$, $C(x,s):=\{y\in\mathbb{R}^{n}:|y^{\prime}-x^{\prime}|<s,|y_{n}-x_{n}|<s\}$, $D(z,s):=\{y\in\mathbb{R}^{n-1}:|y-z|<s\}$. Given an outer measure $\omega$ on $\mathbb{R}^{n}$ and $X\subset\mathbb{R}^{n}$, the trace (restriction) of $\omega$ on $X$ is the outer measure $\omega|X$ on $\mathbb{R}^{n}$ given by $\omega|X(F):=\omega(F\cap X)$. $\mathcal{H}^{n-1}$ denotes the $n-1$-dimensional Hausdorff measure on $\mathbb{R}^{n}$.
@@
       $\mathcal{H}^{n-1}|C(x,s)\cap M=\mathcal{H}^{n-1}|\operatorname{graph}\,u|_{D}$ is a Radon measure due to the area formula [21, Theorem 9.1] . Moreover, one has $\mathcal{H}^{n-1}(\operatorname{graph}\,u|_{D(x^{\prime},s)\setminus G})=\int_{D(x^{\prime},s)\setminus G}|(\nabla u(z),-1)|\operatorname{d}z=0$. ∎
The following preparatory result (3) comprises by the case $M_{0}=\emptyset$ the divergence theorem on bounded open sets with Lipschitz boundary (see also e.g. [21, Remark 9.5] and [1, A8.8]).
@@
Proof. (a) Assume first $\varphi\in C_{c}^{1}(C(x,s))$. By (3) the surface integral is well-defined. Following step one of the proof of [21, Theorem 9.3] the result follows due to $C(x,s)\cap\partial E=C(x,s)\cap M$ by definition (3)(b). Note that the weak gradient $\nabla f_{\delta}$ equals $\frac{1}{2\delta}(-\nabla u(z),1)$ at $z\in G$.
       (b) Now the vector field $\nu_{E}$ is constructed. Applying the result in (a) to all $\varphi\in C_{c}^{1}(C(x,s)\cap C(\bar{x},\bar{s}))$ shows that $\nu$ and $\bar{\nu}$ coincide a.e. on their common domain. There are countably many $C(x_{m},s_{m})$ from the localization in (3), which cover $M$. (Indeed, $M_{0}$ is closed and hence $G_{\delta}$. Since $\partial E$ is compact it follows that $M=\partial E\setminus M_{0}$ is $\sigma$-compact.) Thus we may compose from the corresponding $\nu_{m}$ by means of [21, Theorem 9.1] an $\mathcal{H}^{n-1}$-a.e. determined unit vector field $\nu_{E}$ on $\partial E$ which locally coincides a.e. with $\nu$ from (3).
@@
We are in a position to state and prove our version (3) of the Divergence Theorem. It distinguishes itself in the weakest hypotheses concerning $\partial E$ regularity we know. Thus (3) applies to the proof of the crucial technical result (4).
@@
Indeed, the first claim holds as $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})<\infty$ and $A_{\delta}\downarrow_{\delta}M_{0}$. The very definition of the Minkowski content [14, 3.2.27] implies the second claim.
@@
       According to [21, Theorem 8.1], $\mathcal{H}^{n-1}(D)<\infty$ if $D$ is the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable set $C$ of finite Lebesgue measure. If $C$ is a Lebesgue null set then $\mathcal{H}^{n-1}(D)=0$.
@@
       According to [14, Theorem 3.2.39], $\mathcal{H}^{n-1}(D)=\mathcal{M}^{n-1}(D)=0$ holds, if $D\subset\mathbb{R}^{n}$ is closed and if there is a Lipschitz function $h:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ mapping some bounded $C\subset\mathbb{R}^{n-1}$ onto $D$ or, equivalently, if $h$ is locally Lipschitz defined on $\overline{C}$ with $h(C)=D$. The latter holds because $\overline{C}$ is compact.